A company has liquid assets of $600,000 and current liabilities of $500,000. What is the effect on the quick ratio if the company records an accrual adjustment for salaries of $100,000 and pays accounts payable in the amount of $50,000?

a. The quick ratio will not change as a result of either of these transactions.

b. The accrual adjustment will cause the quick ratio to decrease and the payment of accounts payable will cause an increase in the quick ratio.

c. The accrual adjustment will cause the quick ratio to increase and the payment of accounts payable will not affect the quick ratio.

d. The accrual adjustment and the payment of accounts payable will both cause the quick ratio to decrease.
Answer #1

Present quick ratio = Quick assets/Current liabilities = 600000/500000 = 1.2 Times

Accrual adjustment = 600000/600000 = 1.0 Times

Account payable adjustment = 550000/450000 = 1.22

So answer is b) The accrual adjustment will cause the quick ratio to decrease and the payment of accounts payable will cause an increase in the quick ratio.